# Practitioners’ Perspectives on a Differential Privacy Deployment Registry

### v1

@@ 2.1.5 Accounting
This section describes how privacy loss is accounted for over multiple queries to the underlying data. While this idea is most precisely captured by composition, we also include post-processing. By doing so, we emphasize that post-processed transformations do not incur additional privacy loss (as per properties afforded by DP). This section is required for Tier 3 entries only.
Composition. Composition describes how privacy parameter values accumulate across multiple queries to the underlying data under DP. Examples of composition include sequential composition and parallel composition [13]. Under sequential composition, assuming pure DP, the sum of the privacy parameter value ($\epsilon$) per query is the total privacy parameter value. Parallel composition, again assuming pure DP, says that the total privacy parameter value for a set of queries applied on disjoint subsets of the data is the maximum privacy parameter value used for one of the queries.
Post-processing. Functions applied to the data product after DP is applied. These functions do not incur additional privacy loss. However, the impact of these functions may be important to know about for downstream applications that use the data product. For example, a data curator releasing a data product that includes counts of people in different geographic regions may not want to publish privacy-protected counts that are negative, as they are seemingly nonsensical. Instead, they may set these negative counts to zero. Such transformations are important for downstream data users to be aware of in order to adequately account for uncertainty in their analyses.
@@ 2.1.6 Implementation
Pre-processing, exploratory data analysis, hyperparameter tuning. Prior to applying DP, data curators may query the underlying data to inform choices for the DP deployment—such as how to set the privacy parameters—or otherwise pre-process the data in various ways. Steps such as data cleaning, exploratory data analysis, and hyperparameter tuning (e.g., on privacy parameters) prior to DP being applied all incur privacy loss. These steps and details of whether privacy loss was accounted for, if at all, during these phases should be described here.
Mechanisms. Mechanisms, i.e., algorithms that take the sensitive, unprotected data as input and output a privacy-protected data product. We include both low-level mechanisms (e.g., Laplace or Gaussian [27]) as well as high-level mechanisms, such as the Top Down Algorithm [28]. How these mechanisms were implemented, for example using a particular library, should be described as well. Other security-related implementation details, like what measures were taken to protect against known vulnerabilities, like floating-point and timing channel attacks [29] should also be described here. Last, if code is available for the implementation, it should be linked here.
Justification. Rationale for implementation decisions made by the data curator or others. In short, this information should answer the following questions laid out by Dwork, Kohli, and Mulligan [1]: “What were the assumptions, modelling decisions, thresholds, and subjective decisions made in determining the implementation choices above? Why is the approach a thorough test of the stated assumptions? Was the process validated and verified? If so, how?”
@@ 2.1.7 More information and sources
This section is optional, but recommended for all tiers. It includes public sources for the information described in the rest of the deployment card, if available. Sources might include papers or blog posts. If the data product is publicly available, a link should be included in this section. Finally, any miscellaneous notes about the deployment that do not fit into other sections may be included here.
@@ 3 Differential privacy deployment registry
We instantiated the registry as an interactive interface, built in JavaScript (using libraries such as DataTables.js [^1] https://datatables.net/ and D3.js [^2] https://d3js.org/), HTML, and CSS. We designed the registry to be usable by a practitioner interested in deploying DP or evaluating DP deployments. While the original proposal for a registry [1] suggested its value for a broader range of audiences, like policymakers, we narrow our scope to addressing the needs of practitioners as a first step. As such, we designed the registry keeping the following practitioner-oriented design goals (DGs) in mind.
The registry should help a user:
  • DG1: Understand constituent components of a deployment. In particular, the registry should help a user understand the set of implementation decisions required for a DP deployment and potential implications of these decisions, particularly on privacy.
  • DG2: Find technical & sociotechnical information about a deployment. The registry should help a user find technical implementation decisions about a given deployment, but should also support them in finding information about how these decisions were made, the intended use of the deployment, and other sociotechnical factors. Together, these pieces of information are intended to also support a user in conducting their own analyses or evaluations of deployments.
  • DG3: Identify and compare similar deployments. The registry should help a user identify deployments that are similar to each other in a particular way (e.g., they all employ the same flavor of DP) or meet some criterion (e.g., they were deployed after a certain year). It should also help a user make comparisons across these deployments. For example, the registry should make it easy for a user to learn about the type of deployment model typically employed by a particular data curator, for example.
@@
Figure 2: The registry prototype as an interactive interface. The user has selected the “Broadband Coverage” [30, 31] and “Redistricting Data” [28, 32] deployments.
@@ 3.1 Interface components
Table of deployments. The interactive table (Figure 2B) shows high-level information about each deployment in the registry (DG2). Each row represents a single deployment. The columns in the registry roughly correspond to sections in the deployment card described in Section 2.1, with added columns for tier, description of the data product, and curator. The columns “Implementation” and “Accounting” contain keywords denoting whether more information is available on those topics in the deployment’s card, described below. The column “More info” contains a check mark (✓) if there is more information or sources available in the deployment’s card. To help users identify and compare across similar deployments (DG3), columns are sortable, searchable, and filterable. The table also includes a search bar that applies to the entire table in case users want to search a particular term but do not know which column in which it is likely to appear.
We populated the registry with 21 real-world deployments: ten at Tier 3, eight at Tier 2, and three at Tier 1. We aimed for a mix of deployments across type of curator (large companies, startups, government organizations, etc.) and type of deployment (summary statistics, machine learning model, etc.). Either the first or second author filled out each deployment according to the schema—using publicly-available materials, such as papers and blog posts—and flagged ambiguities along the way. The research team met to resolve ambiguities. Our goal was not to comprehensively document all DP deployments, but rather to curate a substantive enough selection for participants to interact with in our study. We were also limited by time and resources (e.g., how many deployments for which there are detailed enough papers or blog posts to fill out a Tier 3 entry).
Deployment card. To the right of the table is a panel for deployment cards (Figure 2C). Selecting a deployment in the table displays its card. Up to three deployments can be selected at a time, allowing the user to tab between and compare deployments (DG3). The deployment card has sections corresponding to the deployment card schema as described in Section 2.1 (DG1, DG2), with certain sections not appearing for Tier 1 and 2 entries.
@@ 4 Methods
To answer our RQs, we conducted a user study where practitioners interacted with the registry as a technology probe [15], followed by a semi-structured interview about challenges and opportunities of a registry. Given that the registry is not a well-established artifact, our study was exploratory in nature. That is, our goal was to understand the space of possible perceptions surrounding a potential registry. As such, participants were given open-ended tasks which aligned with our exploratory RQs.
@@ 4.1 Participants
We recruited 16 participants through our professional networks. This involved sending our recruitment message to our contacts who work with DP in real-world settings and asking them to share it with others in their network. Prospective participants filled out a brief interest form asking about their prior experience. We invited participants to join our study if they had experience making or evaluating a real-world DP deployment. Upon completing the study, participants received a \$50 gift card. The study was exempted by our university’s IRB.
@@ 4.2 Protocol
The first author led study sessions, each lasting approximately one hour. We began sessions with a brief semi-structured interview to learn about participants’ roles and prior experience with DP. Next, participants completed four open-ended tasks (described below) using the registry. We used a think-aloud protocol [33], and asked follow-up questions along the way. We ended each session by conducting a semi-structured interview about the participant’s experience using the registry as well as challenges and opportunities around the registry gaining adoption. Interview questions are available on OSF. [^3] https://osf.io/73pmk/?view_only=4f48922b022842699084bd3b6f889bed
@@ 4.3 Tasks
Task 1: Examining a known deployment. We first asked participants to familiarize themselves with the registry and its deployments. Then, we asked them to choose one deployment they were familiar with prior to interacting with the registry and read about it in more detail. Next, we asked them to share whether and how their impressions changed, including whether they learned any new technical or sociotechnical details.
Finally, we asked them whether and how their understanding of DP or how it is deployed had changed after interacting with the registry.
Task 2: Evaluating deployments. We asked participants to focus on deployments with which they were not already familiar. We then asked them to choose one deployment that stood out as a “good” deployment (and explain their reasoning) and one deployment that stood out as a “poor” deployment (and explain their reasoning). We prompted them to explain which aspects of deployments they were primarily focusing on, and why.
Task 3: Making a new deployment. For this task, we asked participants to either consider a hypothetical, potential DP use case (details of which are on OSF) or a potential use case they were currently considering in their own work. We asked them to use the registry to determine whether DP is appropriate for the use case and, if so, how they would implement it.
Task 4: Exploring emergent norms. Finally, we asked participants to use the exploratory visualizations or interactive table to reflect on norms or standards that seem to be emerging in the DP ecosystem. We prompted them for their reaction to the standards they identified and asked for their thoughts on how the registry would help or hinder the emergence of such standards.
@@ 4.4 Analysis
We followed a thematic analysis approach [34] to qualitatively analyze participants’ sessions. The first author began by open coding recordings of each participant’s session. Next, they iteratively grouped quotations and observations across participants into themes falling under each RQ. They allowed new themes to emerge throughout this process and followed the method of constant comparison to continually assess the stability of each theme. Then, the entire research team met to iterate on and refine these themes.
@@ 5.2 RQ1: How the registry changes understandings of DP
Participants consistently learned new details about individual deployments that they were already familiar with prior to interacting with the registry (P2, P4, P5, P8, P10–13, P15, P16), particularly after inspecting deployment cards. These details spanned both technical choices—such as details about unprotected quantities, hyperparameter tuning, algorithms, and DP flavors—and sociotechnical information, like rationale for deploying DP or the intended use of a data product. For example, P4 and P11 learned about the specific algorithm used for Apple’s use of DP to protect histograms of emoji usage [5] (Tier 3). As P13 explored the same deployment’s card, they described learning about technical details as well as the purpose of the release:
@@
P10’s learning also focused on sociotechnical aspects. When reading the “Justification” subsection of the deployment card for the U.S. Census Bureau’s redistricting data files [28, 32] (Tier 3), they said they had not previously known that the bureau “ran [an] attack on the previous release and it was this successful,” referring to some of the re-identification rates that prompted the bureau to apply DP for the 2020 Census [32].
@@
For example, when looking at one of Microsoft’s deployments [31, 30], P15 wanted to know more about “this ability to look at devices that are not using Microsoft services” saying they were “intrigued by how that’s possible.” P6 noticed that one of the Wikimedia Foundation’s releases [6] did not collect persistent identifiers for users, saying that it would be useful to understand this process better, because “[i]f your identifier is changing, potentially your privacy loss is increasing,” and they “don’t think this is…a major problem, but potentially it could be if not done with a little bit of care.”
@@
P5 noticed that the intended use of one deployment was to support public health authorities in decision-making during the COVID-19 pandemic, but had follow-up questions about which public health authorities in particular. P4 became curious whether a particular data product, which was used internally at an organization, was deleted after a certain period of time. Together, these examples suggest value in a registry that provides contextual information alongside specific implementation choices, echoing Dwork, Kohli, and Mulligan’s [1] original call.
@@ 5.3.1 RQ2a: Evaluating deployments
Multiple participants (P3, P6, P7, P10–12) explicitly considered whether the deployment would help society and whether applying DP was necessary. As P3 put it, “If you’re using DP for…evil purposes, so what [if] you’re protecting privacy, right?” For them, evaluating a deployment entails an “ethical assessment” around whether it is worthwhile for society to have a particular data product available. P6, P10, and P12 all considered how well-suited DP was to the data. For P10, this meant paying particular attention to the “Justification” subsection of deployment cards to learn whether privacy protection was necessary in the first place. Their evaluation approach entailed first determining whether DP seemed necessary, and only then looking into technical details. Similarly, when identifying poor deployments, P12 looked for deployments where “DP is an overkill” noting that “if an application does not necessarily require such stringent data privacy protection, that could be a bit of a flag for not a great deployment.” P6 further elaborated that certain types of data, like browsing histories, were sensitive enough to warrant DP protections. They also considered whether DP as a technology was a good fit for the type of data release. For example, when evaluating one of the Wikimedia Foundation’s deployments [6], they felt that the data were a good fit for DP since the deployment was comprised of “pageview counts per country…as opposed to highly rich crosstabs.”
@@
A few participants explicitly factored in utility as well (P1, P6, P15, P16). For example, P16 explained the difficulty of judging privacy parameters, saying that “…some privacy [parameter values] would just give garbage utility compared to others,” suggesting that utility also plays a role in their assessment of privacy parameters. Because deployment cards in the registry largely do not contain explicit measures of utility, these participants sometimes made guesses about utility based on factors like the population size of the underlying data. For example, P1 stated that “if [they] see something that has a very strong privacy guarantee and not a huge population, then [they’re] concerned.”
In addition to the privacy unit and privacy parameters, participants considered factors like the deployment model, trust assumptions, and libraries used for implementation. For example, P14 looked not only at privacy loss, but also the deployment model. Specifically, they preferred deployments that limit the number of people with access to the raw data: “…obviously if you can achieve the same $\epsilon$, but somehow you achieve it without having the server or many engineers…inspect the messages, that is great.” They also looked at detailed implementation choices in the deployment cards, noting that they prefer when deployments use well-known DP libraries and open source their code. Along these same lines P5 and P14 both regarded a deployment more highly if it was possible to vet the details shared by the curator. As P5 looked through deployment cards, they specifically wanted to see “Wikipedia style referencing where pretty much every claim has a pointer to where it came from,” in addition to the “More Info and Sources” section at the end of each card.

### v2

@@ 2.1.5 Accounting
This section [^6] An alternative is to group the accounting section within an expanded implementation and analysis section, which would include details about how privacy loss was analyzed (i.e., accounting) and verified (through proofs, stochastic testing, etc.). describes how privacy loss is accounted for over multiple queries to the underlying data. While this idea is most precisely captured by composition, we also include post-processing. By doing so, we emphasize that post-processed transformations do not incur additional privacy loss (as per properties afforded by DP). This section is required for Tier 3 entries only.
Composition. Composition describes how privacy-loss parameter values accumulate across multiple queries to the underlying data under DP. Examples of composition include sequential composition and parallel composition cowan2024hands. Under sequential composition, the total privacy-loss parameter equals the sum of the privacy-loss parameters ($\epsilon$) per query. Under sequential composition, the total privacy-loss parameter ($\epsilon$) equals the sum of the privacy-loss parameters per query. Parallel composition says that the total privacy-loss parameter for a set of queries applied on disjoint subsets of the data is the maximum of the privacy-loss parameter values used for the individual queries.
Post-processing. Functions applied to the data product after DP is applied. These functions do not incur additional privacy loss. However, the impact of these functions may be important to know about for downstream applications that use the data product. For example, a data curator releasing a data product that includes counts of people across geographic regions may not want to publish negative privacy-protected counts, as they are seemingly nonsensical. Instead, they may set these negative counts to zero. Data users should know about these transformations to adequately account for uncertainty in their downstream analyses.
@@ 2.1.6 Implementation
Pre-processing, exploratory data analysis, hyperparameter tuning. Prior to applying DP, data curators may query the underlying data to inform choices for the DP deployment—such as how to set the privacy-loss parameters—or otherwise pre-process the data in various ways. Steps such as data cleaning, exploratory data analysis, and hyperparameter tuning (e.g., on privacy-loss parameters) prior to DP being applied all incur privacy loss. These steps and details of whether privacy loss was accounted for, if at all, during these phases should be described here.
Mechanisms. Mechanisms, i.e., algorithms that take the sensitive, unprotected data as input and output a privacy-protected data product. We include both low-level mechanisms (e.g., Laplace or Gaussian dwork2014algorithmic) as well as high-level mechanisms, such as the Top Down Algorithm abowd20222020. How these mechanisms were implemented (e.g., using a particular library) should be described as well. Other security-related implementation details, like measures taken to protect against known vulnerabilities, such as floating-point and timing channel attacks jin2022we should also be described here. Implementation code should be linked here, if available.
Justification. Rationale for implementation decisions made by the data curator or others. This information should answer the following questions from Dwork, Kohli, and Mulligan dwork2019differential: “What were the assumptions, modelling decisions, thresholds, and subjective decisions made in determining the implementation choices above? Why is the approach a thorough test of the stated assumptions? Was the process validated and verified? If so, how?”
@@ 2.1.7 More information and sources
This section is optional, but recommended across tiers. It includes public sources for the information in the deployment card, if available. Sources might include papers or blog posts. If the data product is publicly available, a link should be included. Finally, any miscellaneous notes about the deployment that do not fit into other sections may be included here.
@@ 3 Differential privacy deployment registry
We instantiated the registry as an interactive interface, built in JavaScript (using libraries such as DataTables.js [^7] https://datatables.net/ and D3.js [^8] https://d3js.org/), HTML, and CSS. We designed to be usable by technically skilled practitioners (e.g., privacy engineers or technical leads) interested in deploying DP or evaluating DP deployments. While the original proposal for a registry dwork2019differential suggested its value for a broader range of audiences, like policymakers, we address the needs of practitioners as a first step. As such, we designed the registry keeping the following practitioner-oriented design goals (DGs) in mind. The registry should help a user:
  • DG1: Understand constituent components of a deployment. The registry should help a user understand the set of implementation decisions required for a DP deployment and potential implications of these decisions, particularly on privacy.
  • DG2: Find technical & sociotechnical information about a deployment. The registry should help a user find technical implementation decisions about a given deployment, as well as information about how these decisions were made, the intended use of the deployment, and other sociotechnical factors. Together, this information is intended to support a user in conducting their own analyses or evaluations of deployments.
  • DG3: Identify and compare similar deployments. The registry should help a user identify deployments that are similar in some way (e.g., used pure DP; deployed after 2020). It should also help a user make comparisons across these deployments. For example, the registry should facilitate learning about the type of deployment model typically employed by a particular data curator.
@@
Figure 2: The registry as an interactive interface. The user has selected deployments by Microsoft pereira2021us, MicrosoftUSBroadbandUsagePercentages and the U.S. Census Bureau abowd20222020, censusHandbook.
@@ 3.1 Interface components
Table of deployments. The interactive table (Figure 2B) shows high-level information about deployments (DG2); each row represents a single deployment. The columns in the registry roughly correspond to sections in the deployment card described in Section 2.1, with added columns for tier, description of the data product, and curator. The columns “Implementation” and “Accounting” contain keywords denoting whether more information on those topics is available in the deployment’s card, described below. The column “More info” contains a check mark (✓) if there is more information or sources available in the card. To help users identify and compare across similar deployments (DG3), columns are sortable, searchable, and filterable. The table also includes a search bar over the entire table so users can search a particular term without knowing which column in which it is likely to appear.
We populated the registry with 21 real-world deployments: ten at Tier 3, eight at Tier 2, and three at Tier 1. We aimed for a mix of deployments across type of curator (large companies, startups, government organizations, etc.) and type of deployment (summary statistics, machine learning model, etc.). Either the first or second author filled out each deployment according to the schema—using publicly-available materials, such as papers and blog posts—and flagged ambiguities along the way. The research team met to resolve ambiguities. Our goal was not to comprehensively document all DP deployments, but rather to curate a substantive enough selection for study participants to interact with. We were also limited by time and resources (e.g., how many deployments for which there are detailed enough papers or blog posts to fill out a Tier 3 entry).
Deployment card. To the right of the table is a panel for deployment cards (Figure 2C). Selecting a deployment in the table displays its card. Up to three deployments can be selected at a time. The user can tab between and compare deployments (DG3). The deployment card has sections corresponding to the schema described in Section 2.1 (DG1, DG2); certain sections do not appear for Tier 1 and 2 entries.
@@ 4 Methods
To answer our RQs, we conducted a user study where practitioners interacted with the registry as a technology probe hutchinson2003technology, followed by a semi-structured interview about challenges and opportunities of a registry. Given that the registry is not a well-established artifact, our study was exploratory in nature. That is, our goal was to understand the space of possible perceptions surrounding a potential registry. As such, participants were given open-ended tasks which aligned with our exploratory RQs.
@@ 4.1 Participants
We recruited 16 participants through our professional networks. We sent our recruitment message to our contacts who apply DP in real-world settings and asked them to share it with their networks. Prospective participants filled out an interest form about their prior experience. We invited participants to our study if they had experience making or evaluating a real-world DP deployment. Upon completing the study, participants received a \$50 gift card. The study was exempted by our university’s IRB.
@@ 4.2 Protocol
The first author led study sessions, each lasting approximately one hour. We began sessions with a brief semi-structured interview to learn about participants’ roles and prior experience with DP. Next, participants completed four open-ended tasks (described below) using the registry. We used a think-aloud protocol wright1991use, and asked follow-up questions along the way. We ended each session by conducting a semi-structured interview about the participant’s experience using the registry as well as challenges and opportunities around the registry gaining adoption. Interview questions are in the appendix.
@@ 4.3 Tasks
Task 1: Examining a known deployment. We asked participants to familiarize themselves with the registry and its deployments. Then, we asked them to choose one deployment they were familiar with prior to interacting with the registry and read about it. Next, we asked them to share whether and how their impressions changed, including whether they learned any new technical or sociotechnical details. Finally, we asked them whether and how their understanding of DP or how it is deployed changed after interacting with the registry.
Task 2: Evaluating deployments. We asked participants to focus on deployments they were not already familiar with. The, we asked them to choose one deployment that stood out as “good” (and explain why) and one deployment that stood out as “poor” (and explain why).
Task 3: Making a new deployment. For this task, we asked participants to either consider a hypothetical DP use case (details of which are in the appendix) or a use case they were currently considering in their own work. We asked them to use the registry to determine whether DP is appropriate for the use case and, if so, how they would implement it.
Task 4: Exploring emergent norms. Finally, we asked participants to use the exploratory visualizations or table to reflect on norms or standards appearing to emerge in the DP ecosystem. We prompted them for their reaction to the standards they identified and asked about how the registry would help or hinder the emergence of such standards.
@@ 4.4 Analysis
We followed a thematic analysis approach braun2006using to qualitatively analyze participants’ sessions. The first author began by open coding recordings of each participant’s session. Next, they iteratively grouped quotations and observations across participants into themes within each RQ. They allowed new themes to emerge throughout this process and followed the method of constant comparison to continually assess the stability of each theme. Then, the entire research team met to iterate on and refine these themes.
@@ 5.2 RQ1: How the registry changes understandings of DP
Participants consistently learned new details about individual deployments that they were already familiar with prior to interacting with the registry (P2, P4, P5, P8, P10–13, P15, P16), particularly after inspecting deployment cards. These details spanned both technical choices—such as details about unprotected quantities, hyperparameter tuning, algorithms, and DP flavors—and sociotechnical information, like rationale for deploying DP or the intended use of a data product. For example, P4 and P11 learned about the specific algorithm used for Apple’s use of DP to protect histograms of emoji usage apple2017 (Tier 3). As P13 explored the same deployment’s card, they described learning about technical details as well as the purpose of the release:
@@
P10’s learning also focused on sociotechnical aspects. When reading the “Justification” subsection of the deployment card for the U.S. Census Bureau’s redistricting data files abowd20222020, censusHandbook (Tier 3), they said they had not previously known that the bureau “ran [an] attack on the previous release and it was this successful,” referring to some of the re-identification rates that prompted the bureau to apply DP for the 2020 Census censusHandbook.
@@
For example, when looking at one of Microsoft’s deployments MicrosoftUSBroadbandUsagePercentages, pereira2021us, P15 wanted to know more about “this ability to look at devices that are not using Microsoft services” saying they were “intrigued by how that’s possible.” P6 noticed that one of the Wikimedia Foundation’s releases adeleye2023publishing did not collect persistent identifiers for users, saying that it would be useful to understand this process better, because “[i]f your identifier is changing, potentially your privacy loss is increasing,” and they “don’t think this is…a major problem, but potentially it could be if not done with a little bit of care.”
@@
P5 noticed that the intended use of one deployment was to support public health authorities in decision-making during the COVID-19 pandemic, but had follow-up questions about which public health authorities in particular. P4 became curious whether a particular data product, which was used internally at an organization, was deleted after a certain period of time. Together, these examples suggest value in a registry that provides contextual information alongside specific implementation choices, echoing Dwork, Kohli, and Mulligan’s dwork2019differential original call.
@@ 5.3.1 RQ2a: Evaluating deployments
Multiple participants (P3, P6, P7, P10–12) explicitly considered whether the deployment would help society and whether applying DP was necessary. As P3 put it, “If you’re using DP for…evil purposes, so what [if] you’re protecting privacy, right?” For them, evaluating a deployment entails an “ethical assessment” around whether it is worthwhile for society to have a particular data product available. P6, P10, and P12 all considered how well-suited DP was to the data. For P10, this meant paying particular attention to the “Justification” subsection of deployment cards to learn whether privacy protection was necessary in the first place. Their evaluation approach entailed first determining whether DP seemed necessary, and only then looking into technical details. Similarly, when identifying poor deployments, P12 looked for deployments where “DP is an overkill” noting that “if an application does not necessarily require such stringent data privacy protection, that could be a bit of a flag for not a great deployment.” P6 further elaborated that certain types of data, like browsing histories, were sensitive enough to warrant DP protections. They also considered whether DP as a technology was a good fit for the type of data release. For example, when evaluating one of the Wikimedia Foundation’s deployments adeleye2023publishing, they felt that the data were a good fit for DP since the deployment was comprised of “pageview counts per country…as opposed to highly rich crosstabs.”
@@
A few participants explicitly factored in utility as well (P1, P6, P15, P16). For example, P16 explained the difficulty of judging privacy-loss parameters, saying that “…some privacy [parameter values] would just give garbage utility compared to others,” suggesting that utility plays a role in their assessment of privacy-loss parameters and that an explicit section for utility may be beneficial to add to the schema. Because deployment cards in the registry largely do not contain explicit measures of utility, these participants sometimes made guesses about utility based on factors like the population size of the underlying data. For example, P1 stated that “if [they] see something that has a very strong privacy guarantee and not a huge population, then [they’re] concerned.”
In addition to the privacy unit and privacy parameters, participants considered factors like the deployment model, trust assumptions, and libraries used for implementation. For example, P14 looked not only at privacy loss, but also the deployment model. Specifically, they preferred deployments that limit the number of people with access to the raw data: “…obviously if you can achieve the same $\epsilon$, but somehow you achieve it without having the server or many engineers…inspect the messages, that is great.” They also looked at detailed implementation choices in the deployment cards, noting that they prefer when deployments use well-known DP libraries and open source their code. Along these same lines P5 and P14 both regarded a deployment more highly if it was possible to vet the details shared by the curator. As P5 looked through deployment cards, they specifically wanted to see “Wikipedia style referencing where pretty much every claim has a pointer to where it came from,” in addition to the “More Info and Sources” section at the end of each card. Such a referencing style may be beneficial to incorporate into a future version of the registry.